OpenStudy (anonymous):

what is the y intercept of the equation 2x+y=7?

OpenStudy (katielong):

well you use the equation: y=?x+? you need to rearrange it so y is at the front, so... 2x+y=7 (minus 2x) y=7-2x (change it) y=-2x+7 now you can see that 7 is last. This means that the y intercept is 7 hope it helps :)

OpenStudy (anonymous):

i think its 7

OpenStudy (anonymous):

2x+y=7 y = -2x + 7 => y intercept = 7